Does the slope of the line having the equation 4x + 2y =3 have a value of 4? why or why not?
Answer by jim_thompson5910(35256)

Solve for y

4x + 2y =3

2y = -4x + 3

y = -4x/2 + 3/2

y = -2x + 3/2

The equation is in the form y = mx+b where

m = -2

b = 3/2

So the slope is actually -2 and NOT 4.